OpenStudy (issachall):

Money is invested into an account earning 4.25% interest compounded annually. If the accumulated value after 18 years will be $25,000, approximately how much money is presently in the account?

OpenStudy (greatlife44):

Let's apply the compound interest formula and solve for the variable which is the principle our starting value. \[A = P(1+\frac{ r }{ n })^{nt}\] \[\frac{ A }{ (1+\frac{ r }{ n })^{nt} } = P \] now we know that the interest is compounded annually so that's once a year. so n = 1 so our formula simplifies to this: \[\frac{ A }{ (1+r )^{t} } = P \] \[\frac{ $25,000 }{ (1+0.0425)^{18} } = P \]

OpenStudy (issachall):

thx